form the top of a tower 100m in height , a ball is is dropped and at the same time another ball is projeceted vertically upward fom the ground with a velocity of 25m/s . find when and where the two ball will meet.?

Grade:Upto college level

1 Answers

Kartik Singh
6 years ago
student-name Ijas Muhammed asked in Physicsfrom the top of a tower 100 m in height a ball is dropped and at the same time a ball is projected upwards from the fround with a velocity of 25m/s . Find when and where the two balls meet? 0 Following 1student-name Aman Dhakal answered this74 helpful votes in Physics, Class XII-ScienceWhen the ball is dropped from top, consider the height from top in which the balls meet each other be "x"For the fall dropped from top, u=0m/s, g=9.8 m/s2, h=x, and time taken be "t",∴x= 0.t + 1/2. (9.8). t2or, x=4.9t2 ..................(1)again, for the ball projected upward, h = (100-x), g=9.8 m/s2, time=t (since both meet at same time), u=25m/s,∴(100-x) = 25.t - 1/2. (9.8). t2or, (100-x) = 25t - 4.9t2or, x = 100 - 25t+4.9t2..............(2)now, euatting (1) and (2), we get,4.9t2 = 100 - 25t+4.9t2or, 4.9t2 = 100 - 25t+4.9t2or, 0 = 100-25.t∴ t = 4 seconds.and, putting t=4 in (1), we get,x = 4.9t2or, x = 78.4 metres.∴ Balls meet each other at the height of 78.4 metres from the top of tower or (100-78.4=)21.6metres above the ground after 4 seconds.